Solve the following elasticity.
1. If price increases from 15 to 25 and the quantity demanded falls from 100 to 80 units, what is the arc elasticity of demand?
2. If price decreases from 40 to 20 and the quantity demanded rises from 25 to 45 units, what is the elasticity demand?
1
Expert's answer
2021-09-28T13:17:21-0400

1. [(80-100)/((80+100)/2)]÷[(25-15)/((25+15)/2)]=(-20/90)÷(10/20)=-400/900=-0.44

2. [(45-25)/25]÷[(20-40)/40)=20/25÷(-20/40)=-40/25=-1.6